# PatchRec: Multi-Grained Patching for Efficient LLM-based Sequential Recommendation

### v1

@@ 4.2.1. Baselines
Figure 5. Performance comparison between PatchRec-I with vanilla with the same item numbers in interaction sequence. The compression ratios of PatchRec-I are 2.27, 3.06, and 2.25 for MovieLens-1M, Goodreads, and MoviLens-100K, respectively. Impressively, with improved efficiency, PatchRec-I does not compromise recommendation accuracy.
@@
Figure 6. Performance comparison between the PatchRec-I with vanilla with comparable numbers of tokens in the interaction sequence. The item number in the interaction sequence is 100 for PatchRec-I, with various compression ratios (cr). For PatchRec-I, the latest-$M$ ($M$=3, 5, 10, 20, 30, 50, 100) items are represented with their textual titles, and earlier items are compressed into item patches. PatchRec-I delivers superior recommendation performance compared to vanilla at the same computational cost.
@@ 4.2.2. Results
Notably, PatchRec-I achieves HR@20 improvements of 9.04%, 26.75%, and 3.45% over vanilla, with compression ratios of 3.44, 6.81, and 3.38 on MovieLens-1M, Goodreads, and MovieLens-100K, respectively. Similarly, PatchRec-S achieves HR@20 improvements of 2.97% and 31.56% on MovieLens-1M and Goodreads, respectively, with higher compression ratios of 9.23 and 13.62. On MovieLens-100K, PatchRec-S achieves comparable performance to vanilla with a compression ratio of 3.85.
These results indicate that both PatchRec-I and PatchRec-S can significantly enhance recommendation performance while drastically reducing the number of LLM input tokens. Overall, the experimental findings highlight the effectiveness of our approach in simultaneously improving recommendation accuracy and reducing computational overhead.
@@ 4.3. Analysis of PatchRec-I
To further assess the performance of PatchRec-I in details, we compare it with vanilla under two distinct experimental settings across three datasets, addressing the following research questions:
@@ 4.3.1. Performance Comparison under the Same Items in the Interaction Sequence (RQ1)
We maintain the same truncation length for interaction sequences, with item counts of 20, 40, 60, 80, and 100, and compare HR@20 of PatchRec-I with vanilla LLM4SR to examine whether the compression of early items in PatchRec-I leads to any loss of recommendation performance. The results are shown in Figure 5.
On the MovieLens-1M and Goodreads datasets, PatchRec-I consistently outperforms vanilla, with relative gains exceeding 5% and 30%, respectively. This improvement may be attributed to the increased information density in the compressed item patches, which mitigates the influence of noisy interactions. On the MovieLens-100K dataset, PatchRec-I achieves performance comparable to vanilla, indicating that the proposed compression strategy does not compromise recommendation accuracy.
Furthermore, under a fixed compression ratio, we observe that as the truncation length (i.e., the items in the interaction sequence) increases across the three datasets, the recommendation performance of PatchRec-I generally improves, unlike vanilla, which doesn’t show steady improvement. This trend suggests that, for PatchRec-I, incorporating more historical interactions enhances the ability to model long-term user behaviors effectively.
(a) Goodreads
(b) MovieLens-100K
Figure 7. The performance-efficiency trade-off of vanilla, PatchRec-I, and PatchRec-S on Goodreads and MovieLens-100K dataset. In the scatter plot, points that are positioned further towards the upper right corner indicate better performance and higher efficiency. both PatchRec-I and PatchRec-S provide a better performance-efficiency trade-off compared to vanilla.
@@ 4.3.2. Performance Comparison under Comparable Numbers of Tokens (RQ2)
We compare the recommendation performance of PatchRec-I with vanilla with comparable token counts for LLM input to evaluate the impact of compression on recommendation performance at an equivalent computational cost. The results are shown in Figure 6.
In all experiments across the three datasets, PatchRec-I processes 100 items, which are represented as item patches or raw textual titles. However, due to varying compression ratios, the number of tokens required differs. For each bar in the figures, both PatchRec-I and vanilla input the same number of tokens. Despite this parity in token usage, PatchRec-I consistently achieves higher HR@20 than vanilla, with relative improvements of at least 7%, 20%, and 2%, and reaching up to 13%, 72%, and 19% on MovieLens-1M, Goodreads, and MovieLens-100K, respectively.
These results demonstrate that across all three datasets, PatchRec-I delivers superior recommendation performance compared to vanilla at the same computational cost, underscoring the efficiency and effectiveness of the proposed compression method.
@@ 4.4. Analysis of PatchRec-S
PatchRec-S further compresses the earliest item patches into session patches, thereby achieving even higher compression ratio. To validate its effectiveness, we compare the compression ratio and HR@20 of three methods — vanilla, PatchRec-I, and PatchRec-S — while maintaining the same number of 100 items in the interaction sequence across different compression ratios on three datasets. The results are shown in Figure 2 and Figure 7.
The figures show that the distribution of PatchRec-I and PatchRec-S points is positioned further towards the upper-right corner compared to vanilla across all three datasets. Specifically:
  • PatchRec-S consistently outperforms vanilla in HR@20 across various compression ratios on the MovieLens-1M and Goodreads datasets. Even at compression ratios as high as 9.23 and 13.62 for MovieLens-1M and Goodreads, respectively, PatchRec-S maintains superior HR@20 compared to vanilla.
These results indicate that both PatchRec-I and PatchRec-S provide a better performance-efficiency trade-off compared to vanilla. Furthermore, the distinction between the two methods demonstrates their complementary strengths: PatchRec-I prioritizes performance, while PatchRec-S emphasizes efficiency.
4.5. Effectiveness of the Patch Pre-training Stage
Table 4. Performance comparison between vanilla, patch pre-training, and data augmentation with dropout on MovieLens-1M dataset. The enhancements achieved by compression are not simply due to the random exclusion of items.
Method HR@10 N@10 HR@20 N@20
Vanilla 0.0933 0.0396 0.1482 0.0491
Dropout 0.0936 0.0402 0.1480 0.0492
Patch Pre-training 0.1026 0.0431 0.1610 0.0532
4.5.1. Impact of the Patch Pre-training Stage
@@ 4.5.2. On Data Augmentation
The experimental results, as summarized in Table 4, reveal that patch pre-training consistently outperforms both the vanilla (i.e., no-compression) and dropout baselines. This demonstrates that the enhancements achieved by our compression strategy are not simply due to the random exclusion of items. Our compression design retains critical patterns while increasing the effective information density of the input. These findings highlight the strength of our method in balancing efficiency and effectiveness, ensuring that compression enhances.
@@ 4.6. Performance and Efficiency for Long-term User Behaviors
Figure 8. Performance (i.e., HR@20) and efficiency (i.e., inference time) comparison of vanilla, PatchRec-I, and PatchRec-S for long sequences (i.e., users with at least 500 behaviors) on MovieLens-1M dataset. PatchRec enhances both performance and efficiency in modeling long-term user behaviors.
@@
To further evaluate the performance and efficiency of PatchRec when applied to even longer sequences, we analyze user groups with more than 500 interactions in MovieLens-1M. This experiment was conducted using a Nvidia A100 GPU with 40 GB of memory. For each test interaction, we maximize the number of historical items visible to all three methods (i.e., vanilla LLM4SR, PatchRec-I, and PatchRec-S) to fully utilize the available GPU memory, which simulates the pragmatism in real-world recommender systems that make full use of all available computing sources and data. Specifically, we truncate the Latest-$K$ ($K$=150, 350) historical items for both vanilla and PatchRec-I, whereas PatchRec-S is capable of accommodating all historical items (1435 at most) within the 40 GB GPU memory constraints.
@@
In the case of PatchRec-I, the latest 20 interactions are retained as textual titles, while earlier interactions are compressed into item patches. For PatchRec-S, the parameter $L$, which denotes the number of items in a session, is set to 20. The experimental results are shown in Figure 8. The figure demonstrates that PatchRec-I achieved a 25.11% improvement in HR@20 with only 12.99% inference time, compared to vanilla. Similarly, PatchRec-S achieved a 20.31% increase in HR@20 relative to vanilla, with a even higher 13.09-fold reduction in inference time. These results indicate that our proposed methods significantly enhance both performance and efficiency in modeling long-term user behaviors, implying that our PatchRec is capable of modeling practical historical behavior lengths via LLM4SR.
5. Future Work
While PatchRec demonstrates promising computational efficiency and recommendation performance, certain limitations remain, which we aim to address in future research:
  • Advanced Compression Designs. Although our proposed compression method is simple yet effective and more suitable for multi-grained patch alignment, there is significant potential for exploring more sophisticated designs. Future work may investigate advanced text-to-item and item-to-session compression methods beyond simple averaging, as well as innovative grouping techniques (i.e., consolidating all items within a specific time window into a single session).
  • Adaptive Multi-Grained Compression. Currently, we employ a uniform, predefined multi-grained compression strategy across all sequences. A promising avenue for future exploration involves developing adaptive approaches that dynamically apply multi-grained compression based on available computational resources. This could include determining, for each sequence, the optimal number of items to compress into session patches, item patches, or retain as raw textual titles.
  • Lifelong Sequence Modeling. While PatchRec enhances efficiency in modeling long-term user behaviors with LLMs, there remains an open challenge in effectively modeling lifelong user histories or infinitely long interaction sequences within constrained computational budgets. Future efforts could focus on innovative architectures or methodologies to address this issue.
By addressing these directions, we aim to further enhance the scalability, flexibility, and effectiveness of PatchRec in real-world recommendation systems.
@@ 6. Conclusion
In this paper, we introduce PatchRec, a simple yet effective compression framework designed for multi-grained modeling of users’ extensive historical behaviors in LLM-based sequential recommendation (LLM4SR). To address the significant challenges posed by LLM constraints and the temporal dynamics of user interactions — mostly untouched in prior works — PatchRec leverages a hierarchical and time-aware compression strategy, enabling a deeper and more nuanced understanding of user preferences. By adopting a two-stage training process, PatchRec effectively adapts LLMs to operate within a multi-grained sequence representation space. Extensive experiments on benchmark datasets demonstrate that PatchRec not only dramatically reduces the input token length for LLMs but also consistently improves recommendation performance.

### v2

@@ 4.1. Experimental Settings
Figure 6. Performance comparison between PatchRec-I with TALLRec with the same item numbers in interaction sequence. The compression ratios of PatchRec-I are 2.27, 3.06, and 2.25 for MovieLens-1M, Goodreads, and MoviLens-100K, respectively. Impressively, with improved efficiency, PatchRec-I does not compromise recommendation accuracy.
@@
Figure 7. Performance comparison between PatchRec-I with TALLRec with comparable numbers of tokens in the interaction sequence. The item number in the interaction sequence is 100 for PatchRec-I, with various compression ratios (cr). For PatchRec-I, the latest-$M$ ($M$=3, 5, 10, 20, 30, 50, 100) items are represented with textual titles, and earlier items are compressed into item patches. PatchRec-I delivers superior recommendation performance compared to TALLRec at the same computational cost.
Implementations. We adopt Llama-3.2-1B-Instruct (Dubey et al., 2024) as the backbone LLM, fine-tuning all its parameters. Each training stage consists of one epoch. To balance computational efficiency and resource allocation, we use a training batch size of 64 for MovieLens datasets and 32 for Goodreads, and a test batch size of 16 for all three datasets. The training employs a cosine learning rate scheduler with a warmup ratio of 0.05, and the learning rates are set to 8e-6,5e-5,1e-5 for MovieLens-1M, Goodreads, and MovieLens-100K, respectively. For each interaction, the latest $K$ historical interactions are selected as the truncated user sequence. All experiments access item titles that appear within their respective datasets to ensure fair evaluation.
Evaluation. To evaluate recommendation performance, we use two key metrics: HitRatio@10/20 and NDCG@10/20. Additionally, to assess efficiency, we introduce the Compression Ratio, defined as the ratio of the number of tokens in the historical item sequence before compression to the number after compression. A higher compression ratio indicates greater efficiency.
4.2. Main Performance Comparison
4.2.1. Baselines
We denote the models trained using patch pre-training and PFT-I as PatchRec-I, and those trained with patch pre-training and PFT-S as PatchRec-S. They are compared against traditional recommender models, efficient long-term sequential recommender models, and LLM-based recommenders.
  • Traditional Recommender Models. We select three sequential recommender models as representatives of different mechanisms for capturing user behavior patterns: GRU4Rec(Hidasi et al., 2016) (RNN-based), Caser(Tang and Wang, 2018) (CNN-based) and SASRec(Kang and McAuley, 2018) (attention-based)  [^6] These models are implemented following (Yang et al., 2023a)..
  • Efficient Long-term Recommender Models. LinRec (Liu et al., 2023) and Mamba4Rec (Liu et al., 2024b) are chosen as representatives of efficient long-term sequential recommendation methods. LinRec leverages linear attention mechanisms, while Mamba4Rec employs Mamba blocks (Gu and Dao, 2023) for efficient modeling of long-term user behaviors.
  • LLM-based Recommenders. We include two representative methods that incorporate LLMs in sequential recommendation: (1) MoRec  [^7] We adopt BERT as the encoder for items’ textual metadata and SASRec as the recommender backbone. (Yuan et al., 2023) enhances traditional recommenders by incorporating textual features encoded by LLMs. (2) TALLRec (Bao et al., 2023b) applies SFT on truncated interaction sequences without compression.  [^8] We adapt TALLRec to the all-ranking task with constrained beam search in item title space, consistent with PatchRec
(a) Goodreads
(b) MovieLens-100K
Figure 8. The performance-efficiency trade-off of TALLRec, PatchRec-I, and PatchRec-S on Goodreads and MovieLens-100K dataset. In the scatter plot, points that are positioned further towards the upper right corner indicate better performance and higher efficiency. Both PatchRec-I and PatchRec-S provide a better performance-efficiency trade-off compared to TALLRec.
@@ 4.2.2. Results
Notably, PatchRec-I achieves HR@20 improvements of 9.04%, 26.75%, and 3.45% over TALLRec, with compression ratios of 3.44, 6.81, and 3.38 on MovieLens-1M, Goodreads, and MovieLens-100K, respectively. Similarly, PatchRec-S achieves HR@20 improvements of 2.97% and 31.56% on MovieLens-1M and Goodreads, respectively, with higher compression ratios of 9.23 and 13.62. On MovieLens-100K, PatchRec-S achieves comparable performance to TALLRec with a compression ratio of 3.85.
These results demonstrate that both PatchRec-I and PatchRec-S significantly reduce the number of LLM input tokens without compromising recommendation performance. Overall, the experimental findings highlight the effectiveness of our approach in simultaneously improving recommendation accuracy and reducing computational overhead.
@@ 4.3. Analysis of PatchRec-I
To further assess the performance of PatchRec-I in details, we compare it with TALLRec under two distinct experimental settings across three datasets, addressing the following research questions:
@@ 4.3.1. Performance Comparison under the Same Items in the Interaction Sequence (RQ1)
We maintain the same truncation length for interaction sequences, with item counts of 20, 40, 60, 80, and 100, and compare HR@20 of PatchRec-I with TALLRec to examine whether the compression of early items in PatchRec-I leads to any loss of recommendation performance. The results are shown in Figure 6.
On the MovieLens-1M and Goodreads datasets, PatchRec-I consistently outperforms TALLRec, with relative gains exceeding 5% and 30%, respectively. This improvement may be attributed to the increased information density in the compressed item patches, which mitigates the influence of noisy interactions. On the MovieLens-100K dataset, PatchRec-I achieves performance comparable to TALLRec, indicating that the proposed compression strategy does not compromise recommendation accuracy.
Furthermore, under a fixed compression ratio, we observe that as the truncation length (i.e., the items in the interaction sequence) increases across the three datasets, the recommendation performance of PatchRec-I generally improves, unlike TALLRec, which doesn’t show steady improvement. This trend suggests that, for PatchRec-I, incorporating more historical interactions enhances the ability to model long-term user behaviors effectively.
@@ 4.3.2. Performance Comparison under Comparable Numbers of Tokens (RQ2)
We compare the recommendation performance of PatchRec-I with TALLRec with comparable token counts for LLM input to evaluate the impact of compression on recommendation performance at an equivalent computational cost. The results are shown in Figure 7.
In all experiments across the three datasets, PatchRec-I processes 100 items, which are represented as item patches or raw textual titles. However, due to varying compression ratios, the number of tokens required differs. For each bar in the figures, both PatchRec-I and TALLRec input the same number of tokens. Despite this parity in token usage, PatchRec-I consistently achieves higher HR@20 than TALLRec, with relative improvements of at least 7%, 20%, and 2%, and reaching up to 13%, 72%, and 19% on MovieLens-1M, Goodreads, and MovieLens-100K, respectively.
These results demonstrate that across all three datasets, PatchRec-I delivers superior recommendation performance compared to TALLRec at the same computational cost, underscoring the efficiency and effectiveness of the proposed compression method.
@@ 4.4. Analysis of PatchRec-S
PatchRec-S further compresses the earliest item patches into session patches, thereby achieving even higher compression ratio. To validate its effectiveness, we compare the compression ratio and HR@20 of three methods — TALLRec, PatchRec-I, and PatchRec-S — while maintaining the same number of 100 items in the interaction sequence across different compression ratios on three datasets. The results are shown in Figure 2 and Figure 8.
The figures show that the distribution of PatchRec-I and PatchRec-S points is positioned further towards the upper-right corner compared to TALLRec across all three datasets. Specifically:
  • PatchRec-S consistently outperforms TALLRec in HR@20 across various compression ratios on the MovieLens-1M and Goodreads datasets. Even at compression ratios as high as 9.23 and 13.62 for MovieLens-1M and Goodreads, respectively, PatchRec-S maintains superior HR@20 compared to TALLRec.
These results indicate that both PatchRec-I and PatchRec-S provide a better performance-efficiency trade-off compared to TALLRec. Furthermore, the distinction between the two methods demonstrates their complementary strengths: PatchRec-I prioritizes performance, while PatchRec-S emphasizes efficiency.
4.5. Effectiveness of the Patch Pre-training Stage
4.5.1. Impact of the Patch Pre-training Stage
@@ 4.5.2. On Data Augmentation
Table 4. Performance comparison between TALLRec, patch pre-training, and data augmentation with dropout on MovieLens-1M dataset. The enhancements achieved by compression are not simply due to the random exclusion of items.
Method HR@10 N@10 HR@20 N@20
TALLRec 0.0933 0.0396 0.1482 0.0491
Dropout 0.0936 0.0402 0.1480 0.0492
Patch Pre-training 0.1026 0.0431 0.1610 0.0532
@@
The experimental results, as summarized in Table 4, reveal that patch pre-training consistently outperforms both the TALLRec (i.e., no-compression) and dropout baselines. This demonstrates that the enhancements achieved by our compression strategy are not simply due to the random exclusion of items. Our compression design retains critical patterns while increasing the effective information density of the input. These findings highlight the strength of our method in balancing efficiency and effectiveness, ensuring that compression enhances.
@@ 4.6. Performance and Efficiency for Long-term User Behaviors
Figure 9. Performance (i.e., HR@20) and efficiency (i.e., inference time) comparison of TALLRec, PatchRec-I, and PatchRec-S for long sequences (i.e., users with at least 500 behaviors) on MovieLens-1M. PatchRec enhances both performance and efficiency in modeling long-term user behaviors.
@@
To further evaluate the performance and efficiency of PatchRec when applied to even longer sequences, we analyze user groups with more than 500 interactions in MovieLens-1M. This experiment was conducted using a Nvidia A100 GPU with 40 GB of memory. For each test interaction, we maximize the number of historical items visible to all three methods (i.e., TALLRec, PatchRec-I, and PatchRec-S) to fully utilize the available GPU memory, which simulates the pragmatism in real-world recommender systems that make full use of all available computing sources and data. Specifically, we truncate the Latest-$K$ ($K$=150, 350) historical items for both TALLRec and PatchRec-I, whereas PatchRec-S is capable of accommodating all historical items (1435 at most) within the 40 GB GPU memory constraints.
@@
In the case of PatchRec-I, the latest 20 interactions are retained as textual titles, while earlier interactions are compressed into item patches. For PatchRec-S, the parameter $L$, which denotes the number of items in a session, is set to 20. The experimental results are shown in Figure 9. The figure demonstrates that PatchRec-I achieved a 25.11% improvement in HR@20 with only 12.99% inference time, compared to TALLRec. Similarly, PatchRec-S achieved a 20.31% increase in HR@20 relative to TALLRec, with a even higher 13.09-fold reduction in inference time. These results indicate that our proposed methods significantly enhance both performance and efficiency in modeling long-term user behaviors, implying that our PatchRec is capable of modeling practical historical behavior lengths via LLM4Rec.
5. Related Work
5.1. LLM for Recommendation
With LLMs demonstrating vast world knowledge and impressive reasoning capabilities across various domains (OpenAI, 2023; Google, 2023; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI, 2024; Shao et al., 2024), their potential for recommendation tasks has garnered significant attention from researchers (Hou et al., 2024b; Yue et al., 2023). Existing work on LLM4Rec can be roughly classified into two categories: (1) LLMs as enhancers: LLMs augment traditional recommender models by processing textual features associated with users and items (Yuan et al., 2023; Hou et al., 2022), and (2) LLMs as recommenders: LLMs directly generate recommendations for users’ next interactions (Gao et al., 2023; Bao et al., 2023b; Rajput et al., 2023; Geng et al., 2022).
The primary focus of this paper is on the latter — LLMs as recommenders — where the core challenge is converting user-item interaction sequences into a format suitable for language modeling. Current item indexing approaches in LLM4Rec primarily fall into two types: (1) Text-based representations, where items are indexed using textual information, such as item titles (Bao et al., 2023b, a), descriptions (Li et al., 2023), or learned semantic IDs (Jin et al., 2024; Hua et al., 2023); (2) Collaborative-incorporated representations, which consider collaborative signals (Zhang et al., 2025; Liao et al., 2024b) in item embeddings.
For both aforementioned types, each item is typically represented by multiple tokens. Due to the quadratic computational complexity $O(n^{2})$ of self-attention mechanism in LLMs w.r.t. sequence length  [^9] The computational complexity of vanilla attention mechanisms is $O(n^{2})$. Recent studies proposed methods to reduce it to approximately $O(n)$, using multiple tokens to represent one item in a user’s interaction sequence becomes computationally demanding. This motivates us to explore more efficient item and even sequence representations, aiming to improve the performance-efficiency trade-off.
5.2. Long Sequence Modeling in LLM4Rec
Recent efforts have made significant strides in addressing the context length limitations of LLMs (Rae et al., [n. d.]; Liu et al., 2024a; Zaheer et al., 2020; Su et al., 2024; Ding et al., 2024). One prominent line of work focuses on context compression, wherein older or less critical information is compressed into a more compact representation. Early method (Chevalier et al., 2023) trains LLMs to summarize or encode lengthy preceding text into compact vectors that can be prefixed to the prompt. Similarly, ICAE (Ge et al., [n. d.]) introduces an in-context autoencoder that encodes long contexts and then reconstructs necessary details when needed. More recent studies (Mu et al., 2023; Yoon et al., 2024) explore prompt compression, demonstrating that selectively retaining only task-relevant facts can preserve performance while substantially reducing input length. Broadly, these compression-based techniques aim to prevent earlier information from overwhelming the model’s capacity as context grows. In contrast to these approaches, PatchRec learns concise embeddings at multiple granularities, enabling efficient modeling of extended histories without relying on explicit textual compression or summarization.
The integration of LLMs into recommendation systems introduces further challenges, particularly the restriction on modeling long user behavior sequences due to limited context length and computational resource constraints (Hou et al., 2024a; Wang et al., 2025, 2019; He et al., 2023). One solution (Bang and Song, 2025; Zhong et al., 2024) enhances LLMs with external long-term memory modules that summarize user interaction histories and maintain dynamically updated profiles. Rather than inputting every past interaction, MemoryBank (Zhong et al., 2024) maintains a running summary that the model can consult when needed. PURE (Bang and Song, 2025) introduces a “Review Extractor” and “Profile Updater” continuously distill evolving user histories into a compact form. Another notable approach, ReLLA (Lin et al., 2024) tackles the challenge by retrieving the top-$K$ interacted items from a user’s history based on semantic relevance to the target item. Distinct from these memory- and retrieval-based methods, PatchRec directly trains LLMs to internalize long-term historical signals through aggregated embeddings, eliminating the need for explicit summarization or retrieval mechanisms.
6. Limitation
While PatchRec shows a strong performance-efficiency trade-off, several limitations remain, especially for real-world deployment:
  • Inference Latency. Despite compression, reliance on LLMs still incurs high inference costs, limiting real-time applicability. Future work can explore model distillation, quantization, or early-exit strategies to reduce latency.
  • Compression Design. PatchRec is simple and effective, but more expressive designs — such as advanced patching or innovative grouping techniques — could be further investigated.
  • Adaptive Granularity. We currently use a fixed multi-grained compression scheme. Dynamically adjusting granularity based on resource constraints or user patterns is a promising direction.
Addressing these limitations will be key to making PatchRec more practical and scalable for real-world recommendation systems.
@@ 7. Conclusion
In this paper, we introduce PatchRec, a simple yet effective compression framework designed for multi-grained modeling of users’ extensive historical behaviors in LLM-based recommendation (LLM4Rec). To address the significant challenges posed by LLM constraints and the temporal dynamics of user interactions — mostly untouched in prior works — PatchRec leverages a hierarchical and time-aware compression strategy, enabling a deeper and more nuanced understanding of user preferences. By adopting a two-stage training process, PatchRec effectively adapts LLMs to operate within a multi-grained sequence representation space. Extensive experiments on benchmark datasets demonstrate that PatchRec not only dramatically reduces the input token length for LLMs but also consistently improves recommendation performance.
@@
Acknowledgements.
This research is supported by the National Natural Science Foundation of China (92270114, U24B20180, 62121002) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).